OpenStudy (caozeyuan):

The continuous random variable X has distribution function given by F(x)=1-e^(-1/2 * x) Find the probability that 2 lies between X and 4X, FInd also the expected value of the width of the interval ( X,4X)

OpenStudy (caozeyuan):

\[F(x)=1-e^{-\frac{ 1 }{ 2 }x} (x \ge 0)\]

OpenStudy (anonymous):

Okay, I think the first step would be to find the probability density function (PDF). Are you familiar with how to do that?

OpenStudy (caozeyuan):

No! I was sleeping during my STAT class, that's why I come here

OpenStudy (anonymous):

To find the probability density function, I believe you can simply take the derivative of the the distribution function, F(x).

OpenStudy (anonymous):

So derivative of F(x) = F ' (x) = ...?

OpenStudy (caozeyuan):

WTF is distribution fuction? Is it short for Cumulative Distribution Function?

OpenStudy (anonymous):

Yes, a continuous distribution function (cdf) is what a distribution function is referring to.

OpenStudy (caozeyuan):

well, dF(x)/dt = 1/2*e^(-1/2 * x)

OpenStudy (caozeyuan):

OMG! Where do you find it?! What a wonderful Xmas gift!

OpenStudy (caozeyuan):

BTW, are you studying further math now?

OpenStudy (anonymous):

I typed "Find the probability that 2 lies between X and 4X" into google and it was one of the first things that came up.

OpenStudy (caozeyuan):

LOL! Google knows what I need!

OpenStudy (anonymous):

I have a bachelor's degree in math. I am not formally studying math at the moment, but I do enjoy helping others learn math (and learning with them).

OpenStudy (caozeyuan):

Why the CDF of a poisson distribution is given as 1-P(X=0) Where P(x=0) means the probability of obtaining o when a random variable X follows a poisson distribution of parameter lambda

OpenStudy (anonymous):

Where does it say that the poisson distribution is 1-P(X=0)? What page number or weblink?

OpenStudy (caozeyuan):

@Yttrium , I'm goinf to be killed by this question!

terenzreignz (terenzreignz):

I wonder why you guys are doing maths on Christmas day of all days...

OpenStudy (caozeyuan):

@terenzreignz , both my math and physics teachers gave me a bunch of questions as "Christmas Gift"! WTF!!!!

terenzreignz (terenzreignz):

I'm here, there is no need to tag me. I'm just here to watch you suffer >:)

OpenStudy (caozeyuan):

@Yttrium , you should find solution here; http://papers.xtremepapers.com/CIE/Cambridge%20International%20A%20and%20AS%20Level/Mathematics%20-%20Further%20(9231)/9231_s10_ms_21.pdf It is question 7

OpenStudy (caozeyuan):

I hate you terenz!

terenzreignz (terenzreignz):

You do? But didn't I save your life at least twice a while back? :-?

OpenStudy (caozeyuan):

Yes, you did , but you are not so kind today

terenzreignz (terenzreignz):

And thus you hate me? :( Oh well, then that's not much incentive for me to step up, is it? :D

OpenStudy (caozeyuan):

I mean " Watch me suffer" is not an appropriate Xmas gift, is it?

terenzreignz (terenzreignz):

awww lighten up, it's Christmas... and besides, can't you take a little humour? :3

OpenStudy (caozeyuan):

I'sorry, maybe Chinese does not have abundant humor

terenzreignz (terenzreignz):

Racist >:) LOL JK I'm only half-Chinese. What's the PDF?

OpenStudy (caozeyuan):

all CIE subjects suck

OpenStudy (caozeyuan):

In particular further math

terenzreignz (terenzreignz):

I cannot see your question in that link.

OpenStudy (caozeyuan):

It should be question 7

OpenStudy (caozeyuan):

That pdf is the answer part

terenzreignz (terenzreignz):

Is this the answer key? D:

OpenStudy (caozeyuan):

if you want the question PDF, change ms at the end into qp

terenzreignz (terenzreignz):

ohhh.... When I said PDF, I meant probability density function -__

OpenStudy (caozeyuan):

LOL! that is so confusing!

terenzreignz (terenzreignz):

Anyway, as stated, your probability density function is the derivative of your distribution function... though, in this question, it's not necessary to differentiate...

OpenStudy (caozeyuan):

I don't understand the CIE MS, it's just a mess

terenzreignz (terenzreignz):

Listen *to ME* >:) The only reason we get the PDF is so that we can integrate it to find probabilities... but if we're going to integrate the PDF (which is the derivative of your distribution function F(x) ) We'll arrive back at F(x) anyway, so let's skip that part. Understand?

OpenStudy (caozeyuan):

I'm still a little bit confused on the concept of CDF, what does it represent?

terenzreignz (terenzreignz):

If F(x) is the CDF, it means F(c) is the probability that x < c

OpenStudy (caozeyuan):

My stat's teacher's lecture drive me to sleep, so I did not hear any words he said.......

OpenStudy (caozeyuan):

ok, that makes sense

OpenStudy (caozeyuan):

so PDF is the probability of x=c, right?

terenzreignz (terenzreignz):

Such an un-dilligent student >:) ...LOL moving on. So, if we let f(x) = F'(x) be the pdf, then the probability that a < x < b is given simply by \[\Large \int\limits_{a}^bf(x) dx=\int\limits_a^bF'(x)dx = F(b)-F(a)\] right?

OpenStudy (caozeyuan):

Yes, I remember that from my Pure Class when my teacher talked about Fundamental Theorem of Calculus

terenzreignz (terenzreignz):

Well then, what we want to find is this: \[\Large x < 2 < 4x\] so this looks... hard, doesn't it?

terenzreignz (terenzreignz):

Because what we DO know how to solve is something of the form a < x < b But not if x is in the boundaries... so what to do?

OpenStudy (caozeyuan):

I don'tknow!

terenzreignz (terenzreignz):

well, shall we just sit here and wait for a miracle? XD

OpenStudy (caozeyuan):

Maybe we should create one

terenzreignz (terenzreignz):

Sure... go ahead, make a miracle happen :D

OpenStudy (caozeyuan):

I can't , I'm not magician

terenzreignz (terenzreignz):

Me neither... so what we do? D:

OpenStudy (caozeyuan):

Mark Scheme is Powerful!

terenzreignz (terenzreignz):

Who/what is a Mark Scheme?

OpenStudy (caozeyuan):

But it does not help when I cannot understand what it is talking about

terenzreignz (terenzreignz):

LOL You mean the answer key? I call it spoilers... it takes the fun out of stuff. And anyway, I already have worked my magic, I was just hoping you'd be able to do it yourself without my interference :D

OpenStudy (caozeyuan):

I'm pretty sure we will have some superposition, as if we are waves, LOL!

terenzreignz (terenzreignz):

Silence. >:(

OpenStudy (caozeyuan):

Let's find a spoiler then

terenzreignz (terenzreignz):

Now, pupil, we have \[\Large x < 2 < 4x\] Suppose we divide everything by x (which is a positive random value) \[\Large 1 < \frac2x < 4\] And now, get the reciprocal of all sides... (which will invert the inequalities) \[\Large 1 > \frac{x}2>\frac14\]And finally, multiply everything by 2, we get \[\Large \frac12 < x < 2\] Much like last time, you're sorely lacking in creativity.... you disappoint me >:( (LOL ^_^)

OpenStudy (caozeyuan):

but the question says \[x \ge 0\] I don't think we should ignore the case when x=0

terenzreignz (terenzreignz):

Of course we shouldn't. But x = 0 already doesn't fit the criterion because, clearly, when x = 0 then it is NOT true that x < 2 < 4x

OpenStudy (caozeyuan):

OMG! you're the Magician!

OpenStudy (caozeyuan):

A spoiler tells me that E(x)=2, but why?

terenzreignz (terenzreignz):

From now on, you will address me as MASTER Terenz >:) LOL Now, all you have to do is integrate the pdf from 0.5 to 2. \[\Large \int\limits_{\frac12}^2f(x)dx = F(2) -F\left(\frac12\right) \]

OpenStudy (caozeyuan):

Ok, that's kindergarden stuff

terenzreignz (terenzreignz):

Show-off >:( Anyway, to get the expectations, this time, you DO need the pdf in its explicit form, so go ahead, differentiate the distribution function F(x) F'(x) = ?

OpenStudy (caozeyuan):

it is\[\frac{ 1 }{ 2}\times e ^{-\frac{ 1 }{ 2 }x}\]

terenzreignz (terenzreignz):

Okay, now, \[\Large f(x) = \frac12e^{-\frac12x}\] Find \[\Large \mathbb{E}[x]=\int\limits_0^\infty xf(x) dx\]

OpenStudy (caozeyuan):

Should we use "by parts" this time?

terenzreignz (terenzreignz):

I thought this was kindergarten stuff, pupil? >:)

OpenStudy (caozeyuan):

No, it's high school stuff

terenzreignz (terenzreignz):

Then it's easy stuff. Integrate it in the manner that seems most appealing to you... (but make sure it's correct :D )

OpenStudy (shamil98):

for a second there i thought you were talking bout the .pdf file, then i scrolled up to learn a new math term "probability density function " lol

terenzreignz (terenzreignz):

We learn something everyday... even on Christmas day :D

OpenStudy (shamil98):

My Christmas present to myself is catching up on missing assignments ( though i don't celebrate it ). What a wonderful gift, right? :D

terenzreignz (terenzreignz):

Time will tell :D

OpenStudy (shamil98):

lol

OpenStudy (caozeyuan):

OK, now I see E(x) is 2

OpenStudy (caozeyuan):

but by parts is so tiring, is there a easy way

terenzreignz (terenzreignz):

Nope... are you telling me you are afraid of working hard, my pupil? D:

OpenStudy (caozeyuan):

I'm afraid of Integration By Parts!

OpenStudy (yttrium):

I was knocked down. -____-"

terenzreignz (terenzreignz):

Conquer your fears. Now, a property of E(x) is that it's linear.. IE E(x + y) = E(x) + E(y) and E(cx) = cE(x) for some constant c. NOW What is the length of the interval [x,4x] ?

OpenStudy (caozeyuan):

what does the term " width of the interval" mean, is it just 3X

terenzreignz (terenzreignz):

Yes it is. Now you need to find the expected value of the LENGTH. Which is 3x. So...? E(3x) = ?

terenzreignz (terenzreignz):

oh sorry, width, not length.

OpenStudy (caozeyuan):

ok, now I see

terenzreignz (terenzreignz):

So... everything is settled, then?